A nurse has a solution that is 25% alcohol and another that is 50% alcohol. How much if the 25% alcohol must she use to make 20 L of a solution that is 40% alcohol?
Answer by stanbon(75887)

A nurse has a solution that is 25% alcohol and another that is 50% alcohol. How much of the 25% alcohol must she use to make 20 L of a solution that is 40% alcohol?
---------------------------
Two-variable solution:
Quantity Eq: t + f = 20 L
Alcohol Eq:0.25t+0.50f = 0.40*20
--------------------------------------
Multiply thru Quantity by 50
Multiply thru Alcohol by 100
--------------------------------------
50t + 50f = 50*20
25t + 50f = 40*20
------------------------
Subtract and solve for "t":
25t = 10*20
t = 8 L (amt. of 25% solution needed)
==================
Cheers,
Stan H.
============